OpenStudy (caitlinnr14):

Solve: 2(x + 6) − (x + 1) = 16

OpenStudy (anonymous):

What do you think you have to do first with this problem?

OpenStudy (caitlinnr14):

hmmmm im stuck..completely on this one please help

OpenStudy (anonymous):

will since we have () we need to get rid of them. so if we distribute 2 into (x+6) What do you get?

OpenStudy (caitlinnr14):

2x+12

OpenStudy (anonymous):

alright now we have 2x+ 12 - (x + 1) we have to get rid of the () so we have to do the same thing but with -1 what do you get?

OpenStudy (caitlinnr14):

-1+-1

OpenStudy (anonymous):

if you distribute -1(x+1) you should get -x + -1 or -x - 1 but now we have \[2x + 12 -x -1 = 16\] now that we know this we can combine like terms and get x by itself for the answer. Do you know how to take it from here?

OpenStudy (caitlinnr14):

ummm

OpenStudy (anonymous):

here i will do it step by step for you :)

OpenStudy (caitlinnr14):

thx

OpenStudy (anonymous):

what we have:\[2x + 12 - x -1 =16\]\[2x - x + 12 - 1 =16\]adding and subtracting like terms\[x + 11 = 16\]subtracting by 11 to get x by itself\[x = 5\] is our final answer

OpenStudy (caitlinnr14):

ohhhhh thank you for ur help

OpenStudy (anonymous):

no problem :)